calculate the minimum number of lines that a grating must have to resolve sodium doublet(5890A and 5896A) in third order.



1
Expert's answer
2021-03-18T15:35:21-0400

Resolving power of grating is given by:

"\frac{\lambda }{\Delta \lambda }=nN"

"n=order\:of\:grating"

"N=number\:of\:lines\:of\:grating"

"Here\:n=3\:\:\:"


"N=\frac{\lambda }{3\Delta \lambda }"


"\Delta \lambda =5896-5890=6"


"N=\frac{\left(\frac{5896+5890}{2}\right)}{3\times 6}"


"N=\frac{5893}{18}"


"N=327"

Minimum number of lines"=" 327